OpenStudy (anonymous):

A teacher has a set of 12 problems to use on a math exam. The teacher makes different versions of the exam by putting 10 questions on each exam. How many different exams can the teacher make? A.12! / 11! B. 66 C. 132 D. 12!

OpenStudy (anonymous):

I think it's B, 66 12!/(10!*2!)=12*11/2!=12*11/2=6*11=66

OpenStudy (anonymous):

http://mathworld.wolfram.com/Combination.html He combines 10 questions of a set of 12.

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

You're welcome.